Lifespan of birds is described using a continuous probability model. The probability that they do not live past x years is expressed as p(x)-0.5| 1-exp|-- -·-| | + 0.311-expl__ 32 If observations in the past that the birds live past 5 years, what is the probability that they will survive between 3 and 6 years?
Answer #1

for above probability of living past 5 years =P(X>5)=1-P(X<5)

=1-(0.5*(1-exp(-5/4)+0.2*(1-exp(-52/32))+0.3*(1-exp(-5/12))=0.5379

probability that survive between 5 and 6 years =P(5<X<6)=P(X<6)-P(X<5)=0.5150-0.4621=0.0529

hence P(survive between 3 and 6 years given survive past 5 years)=P(3<X<6|X>5)

=P(5<X<6)/P(X>5)=0.0529/0.5379=0.0984